A line segment has end points A (-12, -11) and B (6, 19). Find the three points that divide AB into four equal pieces. A) show that the line segment with end points A (5,3) and B (5 ,-5) is a chord of the circle X to the power of 2 + Y to the power of two = 34


1
Expert's answer
2022-03-23T16:59:20-0400

To determine the three points that divide AB into 4 equal parts, we shall find a point X on AB such that X is the midpoint of AB. Then we shall find a point Y on AX such that Y is the midpoint of AX. Finally, we shall find a point Z such that Z is the midpoint of XB


Midpoint of AB "=" "(x,y)" "=" "(\frac{-12+6}{2}, \frac{-11+19}{2})"

Midpoint of AB "= (x,y) = (-3, 4)"


"X(-3,4)"


Midpoint of AX "= (x,y) = (\frac{-12-3}{2}, \frac{-11+4}{2})"

Midpoint of AX "= (x,y) = (\frac{-15}{2}, \frac{-7}{2})"


"Y(\frac{-15}{2}, \frac{-7}{2})"



Midpoint of BX "= (x,y) = (\frac{-3+6}{2}, \frac{4+19}{2})"

Midpoint of BX "= (x,y) = (\frac{3}{2}, \frac{23}{2})"


"Z(\frac{3}{2}, \frac{23}{2})"




Equation of circle : "x\u00b2+y\u00b2=34"


To show that line AB is a chord of the circle, we shall show that one of the points lie on the circle, and the midpoint of line AB is not the centre of the circle


Let "x=-5, y=3"

"x\u00b2+y\u00b2=(-5)\u00b2+3\u00b2=25+9=34"


Hence, "A(-5,3)" lies on the circle


Midpoint of AB "= (x,y) = (\frac{5+5}{2}, \frac{-5+3}{2})"

Midpoint of AB "= (x,y) = (5, -1)"


From "x\u00b2+y\u00b2=34" , it is evident that the centre of the circle is "(0,0)"


"(0,0) \u2260 (5, -1)"


Hence, line AB is a chord of the circle, "x\u00b2+y\u00b2=34"